OpenStudy (anonymous):

There is a Calculus Implicit problem that i cannot seem to find the next step i have the answer i just dont seem to know how they got the positive and negative values in the answer...would anyone know how to solve it? find dy/dx, given x^2+ xy -y^3 = 7 My work includes: 2x+y+x dy/dx -3y^2 dy/dx= 0 After this im not sure exactly how to simplify..

jimthompson5910 (jim_thompson5910):

The goal is to isolate dy/dx. Think of this as one variable. If it helps, let z = dy/dx so \[\Large 2x+y+x*\color{red}{\frac{dy}{dx}} -3y^2*\color{red}{\frac{dy}{dx}}= 0\] turns into \[\Large 2x+y+x*\color{red}{z} -3y^2*\color{red}{z}= 0\] solve the equation for z. When z is fully isolated, replace z with dy/dx

OpenStudy (anonymous):

Thank You so Much, that really helped.

jimthompson5910 (jim_thompson5910):

I'm glad it makes much more sense now

OpenStudy (anonymous):

Yes i appreciate it